# Variational Pseudo Marginal Methods for Jet Reconstruction in Particle Physics

### v1

@@ 2.1.2 Jet reconstruction during inference
Parent node splitting likelihood reconstruction.
@@
$\vspace{-1mm}\ell(t_{i},\lambda,t_{cut},t_{P}^{i},t_{P})=\begin{cases}\vspace{-1mm}f(t_{i}|\lambda,t_{P}^{i}),&\qquad t_{P}>t_{cut}\\ F_{s}(t_{cut},t_{P}),&\qquad t_{P}\leq t_{cut}\,,\end{cases}$ (2)
where $i\in\{L,R\}$ (note that $t_{P}^{L}=t_{P}$, and $t_{P}^{R}=(\sqrt{t_{P}}-\sqrt{t_{L}})^{2}$) and $t_{cut}$ is the cutoff mass squared scale for the binary splitting process to stop (if $t_{i}\leq t_{cut}$, the corresponding node is a leaf of the binary tree). We introduce the cumulative density function $F_{s}(t_{cut},t_{P})$ for a given generative process to stop, i.e. the probability of having sampled a value of $t_{P}<t_{\text{cut}}$, as
@@
$\mathcal{F}(t_{L},t_{R},$ $\lambda,t_{cut},t_{P})=\frac{1}{4\pi}(1-F_{s}(t_{cut},t_{P}))$
×$\times\ell(t_{L},\lambda,t_{cut},t_{P}^{L})\cdot\ell(t_{R},\lambda,t_{cut},t_{P}^{R})\,.$ (4)
@@
Figure 3: Summary of the Csmc framework: A total of $K$ partial states $\{s_{r}^{k}\}_{k=1}^{K}$ are retained as collections of tree structures encompassing the data set. Each iteration within Algorithm 2 comprises three key stages: (1) resampling partial states based on their importance weights $\{w_{r}^{k}\}_{k=1}^{K}$, (2) proposing an expansion of each partial state to form a new one by linking two trees within the forest, and (3) determining the new weights for these new partial states. The illustration above depicts three samples across a jet consisting of four particles, denoted as $A,B,C,$ and $D.$
@@ 2.2 Approximate Inference
and (ii): learning or optimizing the likelihood in Eq. 5 obtained by marginalizing Eq. 6: $\hat{\lambda}=\underset{\hat{\lambda}\in\lambda}{\operatorname*{arg\,max}}\log P(\mathbf{X}|\lambda)\,.$ Variational Inference (VI) offers an approach to tackle both tasks for complex posterior distributions.
Variational Inference.
@@
In the context of Auto-Encoding Variational Bayes (Aevb, both $Q(\tau,\lambda|\mathbf{X})$ and $P(\lambda,\tau,\mathbf{X})$ are jointly trained [Kingma and Welling, 2013, Rezende et al., 2014]. To approximate the expectation in Eq.7, Monte Carlo samples from $Q(\tau,\lambda|\mathbf{X})$ are averaged, and these samples are reparameterized using a deterministic function of a random variable that is independent of $\tau$.
@@
Combinatorial Sequential Monte Carlo.
Csmc, tailored for phylogenetic tree models, approximates a sequence of increasing probability spaces, ultimately aligning with Eq. 5  [Wang et al., 2015]. Csmc employs sequential importance resampling across $\{r\}_{r=1}^{N-1}$ steps to approximate both the unnormalized target distribution $\pi$ and its normalization constant, denoted as $\|\pi\|$, constituting the numerator and denominator in Eq. 5, by $K$ partial states $\{s_{r}^{k}\}_{k=1}^{K}\in\mathcal{S}_{r}$ to form a distribution (see Wang et al. [2015] or Appendix 7),
@@
Csmc, in contrast to standard Smc techniques, manages a combinatorial set representing the realm of tree topologies alongside the continuous branch lengths—both of which are characteristic features of phylogenies [Wang et al., 2015]. Partial states (Monte Carlo samples) are resampled at each rank $r$, ensuring samples remain in high-probability regions, and importance weights are defined as:
@@
where $q(s_{r}^{k}|s_{r-1}^{a_{r-1}^{k}})$ specifies a proposal distribution and $\nu^{-}$ is an overcounting correction defined in Wang et al. [2015]. Resampled states are then extended via proposal distribution simulations (see Fig. 3). This framework allows for the construction of an unbiased estimate of the marginal likelihood, converging in $L_{2}$ norm:
@@
Expanding on the foundation laid by Csmc,  Moretti et al. [2021] introduces Variational Combinatorial Sequential Monte Carlo (Vcsmc) as an approach to learn distributions over phylogenetic trees. Vcsmc employs Csmc as a means to create an unbiased estimator for the marginal likelihood:
@@
In the same work, Moretti et al. [2021] introduces Nested Combinatorial Sequential Monte Carlo (Ncsmc), an efficient proposal distribution, providing an exact approximation to the intractable locally optimal proposal for Csmc. We provide a review of Ncsmc in Appendix 8. The VI algorithms Vcsmc and Vncsmc that utilize the estimators $\hat{\mathcal{Z}}_{CSMC}$ and $\hat{\mathcal{Z}}_{NCSMC}$ each introduce a structured approximate posterior that exhibits factorization across rank events. Each state, denoted as $s_{r}$, is uniquely characterized by its topology, a collection of trees forming a forest, and the corresponding branch lengths. To facilitate reparameterization, discrete terms are either removed from gradient estimates or transformed into Gumbel-Softmax random variables. This transformation yields a differentiable approximation by utilizing a convex relaxation technique applied to the simplex.
3 Methods
Section 3.1 adapts the Csmc approach to perform inference on jet tree structures. Section 3.2 reformulates Vcsmc for inference on global parameters. Subsection 3.2.1 utilizes Vcsmc methodology to learn parameters as point estimates. Subsection 3.2.2 defines a prior on the model parameters to construct a variational approximation on both global and local parameters. The resulting approach is interpreted as a variational pseudo-marginal method establishing connections between pseudo-marginal methods [Andrieu and Roberts, 2009] and Variational Combinatorial Sequential Monte Carlo [Naesseth et al., 2018, Moretti et al., 2021].
3.1 Inference on Tree Structures
Sequential Monte Carlo (Smc) methods [Naesseth et al., 2019, Chopin and Papaspiliopoulos, 2020] are tailored to sample from a sequence of probability spaces, where the final iteration converges to the target distribution. To adapt Csmc for the Ginkgo model, the splitting likelihood at the final coalescent event should align with the likelihood of a Ginkgo jet which is defined as the product of all splitting likelihoods (see Eq. 4). To achieve this, we reformulate the splitting likelihood to induce a dependence on previous splits, as depicted in the recurrence relation:
~$\tilde{\mathcal{F}}(t_{L},t_{R},$ $\lambda,t_{cut})=\frac{1}{4\pi}\cdot(1-F_{s}(t_{cut},t_{P}))$ (12)
×$\times\ell(t_{L},\lambda,t_{cut},t_{P}^{L})\cdot\ell(t_{R},\lambda,t_{cut},t_{P}^{R})$
×~$\times\tilde{\mathcal{F}}(t_{LL},t_{LR},\lambda,t_{\text{cut }})\tilde{\mathcal{F}}(t_{RL},t_{RR},\lambda,t_{\text{cut }})\,.$
In the above, the pair $i,j\in(L,R)\times(L,R)$ defines $t_{ij}$ as the mass squared of the $j$ child of the current $i$ coalescent node and ~$\tilde{\mathcal{F}}\left(t_{L},t_{R},\lambda,t_{\text{cut }}\right)$ for leaf nodes simply evaluates to 1. Eq. 12 represents the tree (sub-tree) likelihood (with root node having squared mass $t_{P}$) as the recursive product of splitting likelihoods. Each term brings its normalization, and the overall normalization is correctly expressed as the product of individual ones. In practice we use dynamic programming to maintain a running sum of cumulative log probabilities across rank events. The Csmc resampling step illustrated in Fig. 3 is now dependent upon the sub-tree splitting history as opposed to only the most recent splitting likelihood.
@@ 3.1 Inference on Tree Structures
In a slight abuse of notation, the probability $\pi(s_{r}^{k})$ of partial state $s_{r}^{k}$ (recall $r\in\{1,\cdots,N-1\}$ denotes the coalescent event and $k$ denotes the Monte Carlo sample) is defined as the product of the probabilities of all disjoint trees $\zeta_{i}$ in the forest $s$: $\pi(s)=\prod\limits_{\zeta_{i}\in s}\pi(\zeta_{i})\,.$ An illustration of a likelihood for a sub-tree along with the likelihood of a partial state is provided in Fig. 4. The Ncsmc algorithm defined in Moretti et al. [2021] can similarly be adjusted to ensure compatibility with this modified framework. The resulting estimators are unbiased and consistent, for proofs see Wang et al. [2015] and Moretti et al. [2021].
Figure 5: Left: Scatterplot comparing log-conditional likelihood of Vncsmc with $K,M=(256,1)$ vs (a) Greedy Search and (b) Beam Search. Across 100 simulated jets, Vncsmc returns higher likelihood on all 100 cases against Greedy Search and 99 cases against Beam Search. Right: Log-conditional likelihood values for Vcsmc (blue) and Vncsmc (red) with $K=\{8,256\}$ (and $M=1$) samples averaged across 5 random seeds. Vncsmc with $K\geq 8$ explores higher probability spaces than the likelihood returned by the simulator, as depicted by the green trace for reference. Vncsmc achieves convergence in fewer epochs than Vcsmc and yields higher values, all while maintaining lower stochastic gradient noise.
@@ 3.2.1 Maximum Likelihood
$Q_{\phi,\psi}\left(s_{1:R}^{1:K},a_{1:R-1}^{1:K}\right)\coloneqq$ (13)
$\prod\limits_{k=1}^{K}q_{\phi,\psi}(s_{1}^{k})\times\prod\limits_{r=2}^{R}\prod\limits_{k=1}^{K}\left[\frac{w_{r-1}^{a_{r-1}^{k}}}{\sum_{l=1}^{K}w_{r-1}^{l}}\cdot q_{\phi,\psi}\left(s_{r}^{k}|s_{r-1}^{a_{r-1}^{k}}\right)\right]\,.$
@@ 3.2.2 Fully Bayesian Inference Using a Variational Pseudo-Marginal Framework
Fully Bayesian Jet Reconstruction.
Simulators rooted in quantum chromodynamics are frequently calibrated to align with the data. In order to circumvent making parametric modeling assumptions, we propose a unification of the generative model and the inference process. The posterior distribution defined in Eq. 5 does not include a prior on $\lambda$ nor does it include variational parameters to learn the proposal distribution in Eq. 13. We define a log-normal distribution over $\lambda\sim\mathcal{N}(\lambda|\mu,\Sigma)$ so that $\lambda$ can be marginalized along with $\tau$. The target distribution can now be specified as follows:
$P_{\theta}(\tau,\lambda|\mathbf{X})=\frac{P_{\theta}(\mathbf{X}|\tau,\lambda)P_{\theta}(\tau|\lambda)P_{\theta}(\lambda)}{P_{\theta}(\mathbf{X})}\,.$ (14)
The generative model parameters can be defined as $\theta=\{\mu,\Sigma\}$ or as the output of a neural network and the proposal parameters $\phi=\{\tilde{\mu},\tilde{\Sigma}\}$ can be shared or separately trained.
Pseudo-Marginal Framework.
@@
This implies that by employing nearly any approximate inference technique to Eq. 16, the resulting marginal distribution of that approximation will serve as an estimate of the actual target distribution. The pseudo-marginal framework by Andrieu and Roberts [2009] designs a Markov Chain $(\theta^{i},u^{i})$ with Eq. 16 as its target distribution.
@@
Figure 6: Left: Log conditional likelihood returned by Vcsmc compared with the exact MAP clustering returned by the cluster trellis Greenberg et al. [2020], Greedy Search and Beam Search, for a dataset of 100 simulated jets generated with Ginkgo. Even with $K=256$ and $N=12$, Vcsmc closely approximates exact cluster trellis values for specific jets. Center Left: A comparison of the running times, highlighting that Vcsmc on $N=64$ significantly outperforms cluster trellis on $N=12$. As $N$ increases, the cluster trellis quickly becomes impractical due to its exponential complexity. Center Right: Csmc and Ncsmc converge to the exact likelihood as $K$ increases (log scale). Right: Inferred Log-Normal Pseudo-Marginal Distribution $\log\mathcal{N}(\lambda|\mu,\Sigma)$ for the Heavy Resonance jet computed using the Ncsmc algorithm. Contours of the log conditional likelihood with stochastic gradient steps on $\mathcal{L}_{NCSMC}$ to learn parameters $\mu=(\mu_{1},\mu_{2})$ traced in red.

### v2

@@ 2.1.2 Jet reconstruction during inference
2.2 Parent node splitting likelihood reconstruction
@@
$\ell(t_{i},\lambda,t_{cut},t_{P}^{i},t_{P})=\begin{cases}f(t_{i}|\lambda,t_{P}^{i}),&\qquad t_{P}>t_{cut}\\ F_{s}(t_{cut},t_{P}),&\qquad t_{P}\leq t_{cut}\,,\end{cases}$ (2)
where $i\in\{L,R\}$ (note that to fix a degree of freedom, $t_{P}^{L}=t_{P}$, and $t_{P}^{R}=(\sqrt{t_{P}}-\sqrt{t_{L}})^{2}$) and $t_{cut}$ is the cutoff mass squared scale for the binary splitting process to stop (if $t_{i}\leq t_{cut}$, the corresponding node is a leaf of the binary tree). We introduce the cumulative density function $F_{s}(t_{cut},t_{P})$ for a given generative process to stop, i.e. the probability of having sampled a value of $t_{P}<t_{\text{cut}}$, as
@@
$\mathcal{F}(t_{L},t_{R},$ $\lambda,t_{cut},t_{P})=\frac{1}{4\pi}(1-F_{s}(t_{cut},t_{P}))\cdot\ell(t_{L},\lambda,t_{cut},t_{P}^{L})\cdot\ell(t_{R},\lambda,t_{cut},t_{P}^{R})\,.$ (4)
@@
Figure 3: Summary of the Csmc framework: A total of $K$ partial states $\{s_{r}^{k}\}_{k=1}^{K}$ are retained as collections of tree structures encompassing the data set. A partial state is defined as a collection of trees, which start out as singleton particles $A$, $B$, $C$ and $D$. Each iteration within Algorithm 2 comprises three key stages: (1) resampling partial states based on their importance weights $\{w_{r}^{k}\}_{k=1}^{K}$, (2) proposing an expansion of each partial state to form a new one by linking two trees within the forest, and (3) determining the new weights for these new partial states. The illustration above depicts three samples across a jet consisting of observed four particles, denoted as $A,B,C,$ and $D.$
@@ 2.3 Approximate Inference
and (ii): learning or optimizing the likelihood in Eq. 5 obtained by marginalizing Eq. 6: $\hat{\lambda}=\underset{\hat{\lambda}\in\lambda}{\operatorname*{arg\,max}}\log P(\mathbf{X}|\lambda)\,.$ Variational Inference (VI) offers an approach to tackle both tasks for non-trivial posterior distributions.
2.3.2 Variational Inference.
@@
In the context of Auto-Encoding Variational Bayes (Aevb, both $Q(\tau,\lambda|\mathbf{X})$ and $P(\lambda,\tau,\mathbf{X})$ are jointly trained (Kingma & Welling, 2013; Rezende et al., 2014). To approximate the expectation in Eq.7, Monte Carlo samples from $Q(\tau,\lambda|\mathbf{X})$ are averaged, and these samples are reparameterized using a deterministic function of a random variable that is independent of $\tau$.
@@
2.3.3 Combinatorial Sequential Monte Carlo.
Csmc, tailored for phylogenetic tree models, approximates a sequence of increasing probability spaces, ultimately aligning with Eq. 5  (Wang et al., 2015). Csmc employs sequential importance resampling across $\{r\}_{r=1}^{N-1}$ steps to approximate both the unnormalized target distribution $\pi$ and its normalization constant, denoted as $\|\pi\|$, constituting the numerator and denominator in Eq. 5, by $K$ partial states $\{s_{r}^{k}\}_{k=1}^{K}\in\mathcal{S}_{r}$ to form a distribution (see Wang et al. (2015) or Appendix B),
@@
Csmc, in contrast to standard Smc techniques, manages a combinatorial set representing the realm of tree topologies alongside the continuous branch lengths—both of which are characteristic features of phylogenies (Wang et al., 2015). Partial states (Monte Carlo samples) are resampled at each rank $r$, ensuring samples remain in high-probability regions, and importance weights are defined as:
@@
where $q(s_{r}^{k}|s_{r-1}^{a_{r-1}^{k}})$ specifies a proposal distribution, $a_{r-1}^{k}\in\{1,\cdots,K\}$ denote resampled ancestor indices with $\mathbb{P}(a_{r-1}^{k}=i)=\nicefrac{{w_{r-1}^{i}}}{{\sum_{l=1}^{K}w_{r-1}^{l}}}$, and $\nu^{-}$ is an overcounting correction defined in Wang et al. (2015). Resampled states are then extended via proposal distribution simulations (see Fig. 3). This framework allows for the construction of an unbiased estimate of the marginal likelihood, converging in $L_{2}$ norm:
@@
Expanding on the foundation laid by Csmc,  Moretti et al. (2021) introduces Variational Combinatorial Sequential Monte Carlo (Vcsmc) as an approach to learn distributions over phylogenetic trees. Vcsmc employs Csmc as a means to create an unbiased estimator for the marginal likelihood:
@@
In the same work, Moretti et al. (2021) introduces Nested Combinatorial Sequential Monte Carlo (Ncsmc), an efficient proposal distribution, providing an exact approximation to the intractable locally optimal proposal for Csmc. We provide a review of Ncsmc in Appendix C. The VI algorithms Vcsmc and Vncsmc that utilize the estimators $\hat{\mathcal{Z}}_{CSMC}$ and $\hat{\mathcal{Z}}_{NCSMC}$ each introduce a structured approximate posterior that exhibits factorization across rank events. Each state, denoted as $s_{r}$, is uniquely characterized by its topology, a collection of trees forming a forest, and the corresponding branch lengths. To facilitate reparameterization, discrete terms are either removed from gradient estimates or transformed into Gumbel-Softmax random variables.
@@
3 Methods
Section 3.1 adapts the Csmc approach to perform inference on jet tree structures. Section 3.2 reformulates Vcsmc for inference on global parameters. Section 3.2.1 utilizes Vcsmc methodology to learn parameters as point estimates. Section 3.2.2 defines a prior on the model parameters to construct a variational approximation on both global and local parameters. The resulting approach is interpreted as a variational pseudo-marginal method establishing connections between pseudo-marginal methods (Andrieu & Roberts, 2009) and Variational Combinatorial Sequential Monte Carlo (Naesseth et al., 2018; Moretti et al., 2021).
3.1 Inference on Tree Structures
Sequential Monte Carlo (Smc) methods (Naesseth et al., 2019; Chopin & Papaspiliopoulos, 2020) are designed to sample from a sequence of probability spaces, where the final iteration converges to the target distribution. However, adapting Csmc for the Ginkgo model requires a crucial modification: ensuring that the splitting likelihood at the final coalescent event reflects the dependence on the entire sub-tree splitting history, not just the most recent split. This dependence is essential for accurately capturing the recursive structure of the tree. Recall that the splitting likelihood is defined as the product of all splitting likelihoods (see Eq. 4).
To achieve this, we reformulate the splitting likelihood to explicitly account for previous splits, as depicted in the recurrence relation:
~$\tilde{\mathcal{F}}(t_{L},t_{R},\lambda,t_{cut})=\frac{1}{4\pi}\cdot(1-F_{s}(t_{cut},t_{P}))$$\times\ell(t_{L},\lambda,t_{cut},t_{P}^{L})\cdot\ell(t_{R},\lambda,t_{cut},t_{P}^{R})$ (12)
×~$\times\tilde{\mathcal{F}}(t_{LL},t_{LR},\lambda,t_{\text{cut }})\cdot\tilde{\mathcal{F}}(t_{RL},t_{RR},\lambda,t_{\text{cut }})\,.$
In the above, the pair $i,j\in(L,R)\times(L,R)$ defines $t_{ij}$ as the mass squared of the $j$ child of the current $i$ coalescent node and ~$\tilde{\mathcal{F}}\left(t_{L},t_{R},\lambda,t_{\text{cut }}\right)$ for leaf nodes simply evaluates to 1. Eq. 12 represents the tree (sub-tree) likelihood (with root node having squared mass $t_{P}$) as the recursive product of splitting likelihoods. Each term brings its normalization, and the overall normalization is correctly expressed as the product of individual ones. In practice we use dynamic programming to maintain a running sum of cumulative log probabilities across rank events. The Csmc resampling step illustrated in Fig. 3 is now dependent upon the full sub-tree splitting history, reflecting the recursive nature of the coalescent process.
In a slight abuse of notation, the probability $\pi(s_{r}^{k})$ of partial state $s_{r}^{k}$ (recall $r\in\{1,\cdots,N-1\}$ denotes the coalescent event and $k$ denotes the Monte Carlo sample) is defined as the product of the probabilities of all disjoint trees $\zeta_{i}$ in the forest $s$: $\pi(s)=\prod\limits_{\zeta_{i}\in s}\pi(\zeta_{i})\,.$ An illustration of a likelihood for a sub-tree along with the likelihood of a partial state is provided in Fig. 4. The Ncsmc algorithm defined in Moretti et al. (2021) can similarly be adjusted to ensure compatibility with this modified framework. The resulting estimators are unbiased and consistent, for proofs see Wang et al. (2015) and Moretti et al. (2021).
@@ 3.2 Inference on Global Parameters
In Section 3.2.1, we define a variational approximation on topologies, while in Section 3.2.2, we outline fully Bayesian inference using a variational pseudo-marginal framework.
@@ 3.2.1 Maximum Likelihood
$Q_{\phi,\psi}\left(s_{1:R}^{1:K},a_{1:R-1}^{1:K}\right)\coloneqq\prod\limits_{k=1}^{K}q_{\phi,\psi}(s_{1}^{k})\times\prod\limits_{r=2}^{R}\prod\limits_{k=1}^{K}\left[\frac{w_{r-1}^{a_{r-1}^{k}}}{\sum_{l=1}^{K}w_{r-1}^{l}}\cdot q_{\phi,\psi}\left(s_{r}^{k}|s_{r-1}^{a_{r-1}^{k}}\right)\right]\,.$ (13)
@@ 3.2.2 Fully Bayesian Inference Using a Variational Pseudo-Marginal Framework
Simulators rooted in quantum chromodynamics are frequently calibrated to align with the data. However, training a simulator separately from the inference process can lead to inefficiencies. To address this, we propose a modification of the posterior distribution defined in Eq. 5 to include a prior on $\lambda$ along with variational parameters to learn the proposal distribution in Eq. 13. We define a log-normal distribution over $\lambda\sim\log\mathcal{N}(\lambda|\mu,\Sigma)$ so that $\lambda$ can be marginalized along with $\tau$. The target distribution can now be specified as follows:
$P(\tau,\lambda|\mathbf{X})=\frac{P(\mathbf{X}|\tau,\lambda)P(\tau|\lambda)P(\lambda)}{P(\mathbf{X})}\,.$ (14)
The generative model parameters can be defined as $\theta\coloneqq\lambda=\{\mu,\Sigma\}$ or as the output of a neural network and the proposal parameters $\phi=\{\tilde{\mu},\tilde{\Sigma}\}$ used in the variational approximation to Eq. 14 can be shared or separately trained.
@@
This implies that by employing nearly any approximate inference technique to Eq. 16, the resulting marginal distribution of that approximation will serve as an estimate of the actual target distribution. The pseudo-marginal framework by Andrieu & Roberts (2009) designs a Markov Chain $(\theta^{i},u^{i})$ with Eq. 16 as its target distribution.
@@
Figure 5: Left: Scatterplot comparing log-conditional likelihood of Vncsmc with $K,M=(256,1)$ vs Greedy Search and Center: Scatterplot comparing log-conditional likelihood of Vncsmc with $K,M=(256,1)$ vs Beam Search. Across 100 simulated jets, Vncsmc returns higher likelihood on all 100 cases against Greedy Search and 99 cases against Beam Search. Right: Log-conditional likelihood values for Vcsmc (blue) and Vncsmc (red) with $K=\{256\}$ (and $M=1$) samples averaged across 5 random seeds. Vncsmc achieves convergence in fewer epochs than Vcsmc and yields higher values, all while maintaining lower stochastic gradient noise. Additional experiments demonstrating the effect of $K$ appear in Fig. 9 of the Appendix.